OpenStudy (anonymous):

z ÷ 3 – 5z + 4(z – 2) when z = 12

sammixboo (sammixboo):

Plug in the info \(\rm \color{#20bd23}{\dfrac{12}{3}~-~5\times12~+~4(12~-~2)}\) First do what is in the parenthesis... What is 12 - 2?

OpenStudy (anonymous):

10

sammixboo (sammixboo):

Correct, so we have \(\rm \color{#20bd23}{\dfrac{12}{3}~-~5\times12~+~4\times 10}\) Now we divide/multiply starting from the left side of the equation, so what is 12/3?

OpenStudy (anonymous):

4

sammixboo (sammixboo):

Good, so we have \(\rm \color{#20bd23}{4~-~5\times12~+~4\times10}\) Now still dividing/multiplying startinf from the left side of the equation, what is 5 x 12?

OpenStudy (anonymous):

60

sammixboo (sammixboo):

Yes, so we have \(\rm \color{#20bd23}{4~-~60~+~4\times10}\) What is 4 x 10?

OpenStudy (anonymous):

40

sammixboo (sammixboo):

Good, now we have \(\rm \color{#20bd23}{4~-~60~+~40}\) Now subtracting/adding from the left side of the equation, what is 4 - 60?

OpenStudy (anonymous):

56

sammixboo (sammixboo):

Ah try again.. What is 4 - 60 in that order, not 60 - 4

OpenStudy (anonymous):

-56

sammixboo (sammixboo):

Yes, so we have -56 + 40 Can you solve that for me?

OpenStudy (anonymous):

-16

sammixboo (sammixboo):

And that is your answer :)